Instrument Log

 

IRSI C1 Lamont, OK

 

Date

Technician

Comments

6/22/2016

MS

Hatch was stuck open and the Current System Status was displaying Tuesday June 21 17:31:32 GMT. Program had stopped. Power cycled the instrument and manually closed the hatch at 1409 GMT. RTS at 1424 GMT

5/24/2016

MS

Temperature has been above normal in the Enclosure. Upon mentor request the AC was turned on in the enclosure at 1338 GMT. RTS at 1353 GMT.

4/22/16

Mentor

Runtime and airmass calibration files need to be updated. Mentor updated runtime and airmass calibration files while on site at 1639 GMT.

2/1/16

KT

There seems to be an issue with the IR camera showing a grainy effect in the image. At 1720 GMT stopped the acquisition and disconnected the ethernet jack from the IR camera. At 1729 GMT connected the ethernet jack to the IR camera and started acquisition. RTS at 1735 GMT.

12/15/15

MS

Computer needed to be restarted following scheduled software and security patch updates. At 1404 GMT computers were updated and rebooted by the IT department, some data might be lost.

7/20/15

MS

IRSI has been out for repairs since 6/12/2015. IRSI is back as of 7/20/2015. Reinstalled IRSI at 1445 GMT 7/20/2015. RTS at 1515 GMT.

6/12/15

MS

Due to high temperature failure on the IRSI, the Mentor has requested that it be shut down. Shut down IRSI at 1808 GMT.

6/12/15

MS

IRSI program stopped running at 0000 GMT. Restarted IRSI program at 1451 GMT. RTS at 1506 GMT.

6/11/15

MS

IRSI program has stopped running at 1800 GMT. Restarted IRSI program at 1915 GMT. RTS at 1930 GMT.

6/11/15

JEM

The IRSI continues to overheat even with the addition of the RV style air conditioning unit and all other attempts that have been made to properly cool the unit. Ensured that the thermoelectric cooler is getting the proper voltage to function properly and it is. There is no further action that can be taken by site ops at this time to correct the issue. Also reinstalled the seal underneath the unit that the vendor had installed since removing it did not help the enclosure temperatures even though it greatly helped the temperature around the exterior portion of the cooler. Will notify the vendor and mentor that the vendor needs to provide another solution to correct the problem. RTS @1445 GMT.

6/11/15

MS

IRSI program has stopped running at 1900 GMT on 6/10/2015,

Restarted IRSI program at 1418 GMT. RTS at 1433 GMT.

6/10/15

JEM

The IRSI continues to overheat even with the addition of the RV style air conditioning unit and while external temperatures around the cooler are being held below 80F. Removed the ventilator fan from the external fins of the cooler. There was a significant amount of dirt and debris that were blocking airflow from the ventilator through the fins. Thoroughly cleaned the dirt and debris from the fins and reinstalled the ventilator fan @1700 GMT. Will continue to monitor the IRSI to see if this has helped the temperature regulation. RTS @1715 GMT.

6/9/2015

MS

IRSI had stopped collecting data at 1800 GMT. Restarted IRSI at 1939 GMT. RTS at 1954 GMT.

6/9/15

JEM

The IRSI continues to overheat even with the addition of the RV style air conditioning unit. The air conditioning unit is designed to blow cold air on the exterior portion of the thermoelectric cooler underneath the IRSI that is used to cool the interior of the enclosure. The vendor also sealed the other side of the underneath portion of the IRSI. This sealing has caused 2 problems. The first problem is the insulation it too close to the heat sink fins of the cooler not allowing proper air flow. The second problem is the space underneath the IRSI is divided by a steel plate with approximatly a 1" gap at the top and bottom. The air conditioner is on the opposite side of the space from the cooler. This is not allowing for proper air flow and basically the hot air is staying trapped around the cooler underneath the unit. Temperatures around the cooler are constantly exceeding 100F while the other half of the space is relatively cool. Removed the seal on the end of the underneath space to allow better air flow across the cooler fins and allow the air conditioning unit to force the hot air out of the space and away from the cooler. Temperatures around the cooler are consistently staying at or below 80F now. Will leave the seal out to see how the IRSI operates now. RTS @1515 GMT.

6/8/15

JEM

Instrument vendor installed RV style air conditioning unit to the IRSI, but it is not properly wired in to AC power, so that needs to be corrected. There is a din rail mounted power supply mounted under the air conditioning unit. This power supply is mounted too close to the ground, is not designed for outdoor use and will be exposed to large amounts of condensation in its current location. Removed this power supply and mounted it inside the 8"x8"x4" enclosure where the electrical outlets are. Connected this power supply to UPS power and kept the air conditioning unit on raw power. Now in the case of a power bump or outage, the thermostat will still function and will not have to be powered on a reset. Ran the air conditioner power cord and the 12V power cord for the thermostat in conduit between the electrical box and the IRSI. RTS @1930 GMT.

6/8/15

MS

Vendor is on site putting in AC unit for the IRSI at 1400 GMT. Rebooted IRSI program at 1445 GMT. RTS 1500 GMT.

6/3/15

MS

IRSI program has stopped collecting data at 2100 GMT on 6/2/2015. Shutdown IRSI at 1517 GMT on 6/3/2015. Shutdown was requested by Mentor.

6/2/15

MS

Instrument stopped collecting data at 1900 GMT on 6/1/2015.

Rebooted IRSI at 1411 GMT. RTS at 1426 GMT.

6/1/15

MS

System locked up at 2100 GMT on 5/29. Restarted IRSI and RTS at 1357 GMT.

5/29/15

MS

Last image shown is on 5/29/2015 at 00:00 GMT. Unable to Radmin IRSI computer. Power cycled IRSI at 1415 GMT. RTS at 1430 GMT.

5/28/15

PD

Last image shown is on 5/26/15 at 1611 GMT. Unable to Radmin IRSI computer. Power cycled IRSI at 1431 GMT. RTS at 1446 GMT.

5/19/2015

MS

Program stopped on the IRSI at 2100 GMT on 5/18/2015. Restarted IRSI at 1431 GMT on 5/19/2015. RTS at 1446 GMT.

5/18/2015

MS

Program stopped on the IRSI at 1600. Restarted IRSI at 1817 GMT. RTS at 1832 GMT.

5/18/2015

MS

Program stopped running at 2100 GMT 5/16/2015. Reset IRSI program at 1423 and it didn't take, reset a second time at 1527 GMT. RTS at 1542 GMT.

5/5/15

MS

Program stopped at 2100 GMT on 5/4. Restarted program and RTS at 1514 GMT.

5/4/15

MS

Program stopped at 1700 GMT. Restarted program and RTS at 1810 GMT.

5/4/15

MS

Program stopped at 0100 GMT on 5/2. Restarted program and RTS at 1514 GMT.

5/1/15

MS

Program stopped at 1930 GMT. Restarted program and RTS at 2045 GMT.

5/1/15

MS

Program stopped at 2300 GMT on 4/30. Restarted program and RTS at 1415 GMT.

4/30/15

MS

Program stopped at 1900 GMT. Restarted program and RTS at 2014 GMT.

4/30/15

MS

Program stopped at 2300 GMT on 4/29. Restarted program and RTS at 1424 GMT.

4/29/15

MS

Program stopped at 2000 GMT. Restarted program and RTS at 2030 GMT.

4/25/2015

MS

Program had stopped at 2000 GMT on 4/25/15. Restarted program at 1432 GMT. RTS at 1432 GMT

4/22/15

MS

Program stopped running at 2100 GMT on 4/21/2015. Restarted program at 1458 GMT on 4/22/2015, RTS 1513 GMT.

4/17/15

MS

IRSI program was not running at 1350 GMT. Rebooted IRSI at 1435 GMT and it worked for a couple of hours and then shut back off at 1600 GMT. Rebooted again at 1900 GMT. RTS 1915 GMT.

4/13/15

KT

The instrument wasn't online nor was the hatch opening or closing. At 1406 GMT power cycled the instrument computer. At 1417 GMT the instrument came online and was operational. RTS at 1421 GMT.

4/9/2015

MS

Program for the IRSI stopped responding at 2100 GMT on 4/8/2015. At 1431 GMT on 4/9/2015 power cycled the IRSI.

4/8/2015

MS

Program for the IRSI stopped responding at 1800 GMT on 4/7/2015. At 1413 GMT on 4/8/2015 power cycled the IRSI. RTS at 1428 GMT.

4/7/15

MS

Program for the IRSI stopped responding at 0100 GMT on 4/7/2015.

At 1443 GMT on 4/7/2015 power cycled the IRSI. RTS 1443 GMT.

4/3/15

KT

SDS reported no data being collected from the instrument. At 1544 GMT could not communicate with the instrument computer. At 1545 GMT power cycled the instrument computer. RTS at 1559 GMT.

4/2/15

MS

Program for the IRSI stopped responding at 2200 GMT on 4/1/2015. At 1455 GMT on 4/2/2015 power cycled the IRSI. RTS at 1510 GMT.

3/31/2015

MS

Program for the IRSI was not responding at 1505 GMT. Power cycled the IRSI at 1505 and RTS at 1520 GMT.

3/30/2015

MS

Program for the IRSI was not responding at 1505 GMT. Power cycled the IRSI at 1505 and RTS at 1520 GMT.

3/24/15

KT

Program was not responding at 1457 GMT. Power cycled IRSI at 1500 GMT. RTS at 1515 GMT.

3/23/15

MS

Program was not responding at 1500 GMT. Power cycled IRSI at 1515 GMT.

3/17/15

MS

Program was not responding at 1300 GMT. Restarted the IRSI program at 1445 GMT.

3/13/15

MS

Hatch was not responding starting at 3/12/2015 2300 GMT. Power cycled IRSI on 3/13/2015 1435 GMT. IRSI hatch is now responding.

3/12/15

MS

Sun Tracker was not responding at 1513 GMT. Restarted program on the SIRS at 1528 GMT.

2/17/15

CAM/JEM

Hatch blackbody temperature was reading 0.0C. T/S found the sensor wires as well as the heater power wires were completely broken in two inside the braided shield. Repaired cables at 2130 GMT and restarted instrument. Temperature reading returned to normal but no images displayed on the console. Database Error displayed: Table ./solmirus/infraredDATA is marked as crashed and last repair failed. Rebooting computer and power cycling instrument produced no change. Notified mentor.

12/17/14

KT

The hatch was not opening or closing. Could not connect to the instrument computer. At 1512 GMT power cycled the instrument computer. After power cycle was completed I was able to log into the instrument computer and the hatch was opening and closing. System appears to be online and performing normal operations. RTS at 1527 GMT.

12/2/14

KT

Could not connect to the computer. The instrument hatch was not opening or closing. At 1505 GMT rebooted the instrument computer. At 1510 GMT system was online and operational. RTS at 1520 GMT.

9/30/14

KT

The IRSI hatch was not opening or closing. At 1405 GMT power cycled the computer. At 1410 GMT the instrument came online, the hatch was opening and closing, system is operational at this time. RTS at 1420 GMT.

9/29/14

KT

The hatch was not opening or closing. At 1425 GMT trying logging into the computer. The computer was locked up. Power cycled the instrument computer. After instrument computer came online the hatch began opening and closing and I was able to log into the instrument computer. the image display is still not updating. RTS at 1440 GMT.

9/26/14

KT

The hatch was not opening or closing. At 1420 GMT rebooted the computer. At 1424 GMT the hatch began opening and closing. System is online and operational but yet still has yet to produce up to date images. RTS at 1435 GMT.

9/23/14

KT

SDS reported not being able to connect to the instrument. At 1425 GMT during routine maintenance the instrument computer was not running nor was the hatch opening and closing. At 1429 GMT power cycled the instrument computer. At 1435 GMT after the computer performed a full reboot the hatch began opening and closing. Logged into the instrument computer and all appears to operational at this time. RTS at 1440 GMT.

9/22/14

CAM

Steven from Solmirus called and advised he had inadvertently shut the IRSI computer down and requested that the IRSI be power cycled. Power cycled computer at 2217 GMT. RTS status O.K.

9/22/14

KT

SDS reported not being able to connect to the instrument. At 1453 GMT during routine maintenance the instrument computer was not running nor was the hatch opening and closing. At 1455 GMT power cycled the instrument computer. At 1500 GMT after the computer performed a full reboot the hatch began opening and closing. Logged into the instrument computer and all appears to operational at this time. RTS at 1508 GMT.

8/11/14

KT

The hatch was not opening or closing. At 1555 GMT rebooted the computer. The hatch began to open and close. Logged into the computer and began recording the values needed and after 5 minutes the hatch quit opening and closing and the computer program lost connection. At 1630 GMT rebooted the computer once again and the hatch began normal operations. RTS at 1635 GMT.

8/7/14

KT

The hatch was not opening or closing. At 1805 GMT rebooted the computer program. The hatch could manually be opened and closed. After rebooting the program the system came online but hatch would still not function properly. RTS at 1820 GMT. Trouble shooting still in progress. Contacted instrument mentor.

7/28/14

CAM

Steven from Solmirus called and requested assistance from Site Ops. IRSI computer is overheating to the point of shutting down and he wants to stop the program and slow the CPU down until a cooling solution can be implemented. Connected a monitor and keyboard directly to IRSI computer and entered BIOS setup. Changed CPU ratio to 10X 2000 MHz as requested. Saved new setting and RTS at 2100 GMT. Blackbody hatch was closed remotely and program left not running.

7/25/14

KT

Stephen from Solmaris called and requested that the IRSI computer be power cycled so he is able to log in to further trouble shoot on going computer issues. At 1600 GMT power cycled the IRSI computer. Instrument came online and began normal operations allowing access to perform computer trouble shooting. RTS at 1615 GMT.

7/23/14

CAM

Instrument vendor called and requested the instrument be power cycled. Instrument has been down for several days, but Site Ops were previously asked not to restart until notified by vendor. Power cycled instrument at 2202 GMT. Verified instrument was accessible via the web interface.

7/15/14

KT

The hatch was not opening or closing. Could not log into instrument. At 1340 GMT opened up instrument enclosure. Was able to manually open and close hatch. Power cycled the computer system. Hatch began opening and closing on it's own. Went to computer located inside the Optical trailer and was able to log in to the instrument. System is online and operational at this time. RTS at 1355 GMT.

7/14/14

KT

SDS reported that the instrument had not collected for 69 minutes. At 1830 GMT during inspection of the instrument, the same issue that was found this morning seemed to have happened again. The hatch either stopped opening or closing but was found vertical and not moving. Opened the enclosure and power cycled the computer system once again. Checked with David Swank and after power cycling the instrument was now reporting files. RTS at 1845 GMT.

7/14/14

KT

The hatch was not opening or closing. Could not log into instrument. At 1435 GMT opened up instrument enclosure. Was able to manually open and close hatch. Power cycled the computer system. Hatch began opening and closing on it's own. Went to computer located inside the Optical trailer and was able to log in to the instrument. System is online and operational at this time. RTS at 1450 GMT.

6/10/14

KT

The hatch on the instrument was not opening or closing. Instrument vendor arrived on site to troubleshoot. After further inspection, the gear that controls the hatch opening and closing had sheared in half and needed to be replaced. Assisted instrument vendor in replacement of new gear assembly. Instrument is back on line and operational. RTS at 1700 GMT.

5/29/14

CAM

Condensation inside the visible instrument dome. Instrument mentor requested the dome/ring assembly be removed, the dome removed from the ring and then only the dome be reattached over the visible instrument using silicone. Placed instrument in OFF mode at 1742 GMT. Removed and reinstalled dome as per vendor supplied directions. RTS at 1900 GMT.

5/15/14

CAM

Cleaned water spots and dust from the IR lens. Found there is an alignment/clearance issue with the visible dome and its blackbody. The edge of the blackbody is actually rubbing on the dome as it comes down (closes) and has damaged the dome. The blackbody also rubs on the black ring around the base of the dome as it come to rest in the fully closed position. Then when the blackbody opens, it snags on the ring just momentarily, but does go ahead and open. It does not appear to rub the dome when it opens, only when it closes. Notified mentor and vendor.

5/15/14

CAM

Used clamp on ammeter to measure IRSI current draw. Initially pulling 2.0 amps on average, varying from 1.8 to 2.1 when the blackbody arm would move.  After several minutes, current draw jumped to 5.0 amps fairly steady, with the highest ever observed being a brief 5.6A.  Assume the heater or cooler came on, but not sure which.  Ambient temperature was around 58F at time of measurements. After a few more minutes, current draw dropped back to around 2.0 amps. Notified mentor of findings.

5/7/14

Vic Morris

Turned on blackbody heater switch at 1545 GMT to allow calibration by vendor. Heater is controlled via software so vendor will turn it on as necessary when ready to do calibrations. Switch will be left on until notified by vendor or mentor to turn it off.

5/2/14

Vic Morris

Solmirus ASIVA (IRSI) instrument installed by mentor at 1826 GMT.